# Large Language Models for Travel Behavior Prediction

### 30 Nov 2023 v1

@@ Abstract
Travel behavior prediction is a fundamental task in transportation demand management. The conventional methods for travel behavior prediction rely on numerical data to construct mathematical models and calibrate model parameters to represent human preferences. Recent advancement in large language models (LLMs) has shown great reasoning abilities to solve complex problems. In this study, we propose to use LLMs to predict travel behavior with prompt engineering without data-based parameter learning. Specifically, we carefully design our prompts that include 1) task description, 2) travel characteristics, 3) individual attributes, and 4) guides of thinking with domain knowledge, and ask the LLMs to predict an individual’s travel behavior and explain the results. We select the travel mode choice task as a case study. Results show that, though no training samples are provided, LLM-based predictions have competitive accuracy and F1-score as canonical supervised learning methods such as multinomial logit, random forest, and neural networks. LLMs can also output reasons that support their prediction. However, though in most of the cases, the output explanations are reasonable, we still observe cases that violate logic or with hallucinations.
@@ 1 Introduction
Understanding human travel behavior has always been a crucial part of transportation planning. Travel behavior is a broad concept that consists of many attributes, such as the mode of transportation, the purpose of the travel, the choice of destination, and the time of departure. Consider travel mode choice as an example, which refers to how people make decisions on their means of travel. In the task of predicting mode choices, making accurate predictions is useful for both agencies and individuals. For agencies, it facilitates providing better transit services and forecasting travel conditions in the short term, while informing transportation infrastructure investments, better policies, and land use planning in the long run. In the traditional Four-Step Model for future demand and performance analysis, the mode share of a given population is estimated using mode split models (Ben-Akiva and Lerman, 1985). For individuals, on the other hand, mode choice prediction embedded in map applications helps them to make sound travel decisions. Conventional mode choice analysis relies on using numerical data to construct a mathematical model to represent human preferences (Mo et al., 2018, 2021). One example is the discrete choice model, where mode choice probabilities are calculated assuming people make choices to maximize their utilities (Ben-Akiva and Lerman, 1985).
@@
Since last year, large language models (LLMs) have generated a tremendous amount of excitement in research due to their advanced language interaction and interpretation capabilities exhibited by-products such as ChatGPT. Having been trained on extremely large corpus and with parameters to the size of more human hundreds of billions, LLMs are able to interpret and execute diverse types of intricate tasks with a certain level of intelligence (Zhao et al., 2023). In transportation, studies have shown the abilities of GPT in making mobility predictions which outperforms state-of-the-art time-series machine learning models (Wang et al., 2023b). However, the question of whether LLMs can tackle the problem of travel behavior predictions is unclear.
In response to the above inquiry, this research aims to analyze the ability of LLMs to make travel behavior predictions. Specifically, this research focuses on travel mode choice prediction as a case study. Through prompt engineering, we are able to establish a zero-shot prompting framework that describes the task of the travel mode choice prediction, travel characteristics, and passenger attributes, as well as guiding the LLM with transportation domain knowledge. We demonstrate that although LLM-based predictions do not require any specific data inputs, they can achieve competitive performance than the classical models such as multinomial logit, random forest, and neural networks.
This paper makes the following contributions:
  • 1.
    This paper presents a framework to make travel behavior predictions with LLMs. To the best of the knowledge of the authors, this is the first effort to utilize the semantic power of LLMs in travel behavior prediction.
  • 2.
    The research demonstrates the accuracy and robustness of the framework by testing and evaluating it with real-world survey data and comparing the performance with classical benchmark models.
  • 3.
    The research proposes a paradigm shift in travel forecasting from using numerical data to establish mathematical models, to leveraging the reasoning ability of LLMs to predict travel behavior without training samples.
@@ 2.1 Recent developments in large language models
Large Language Models (LLMs) typically refer to the language models based on the Transformer architecture with more than hundreds of billions of parameters, which have been shown to have abilities to excel in many complex tasks (Zhao et al., 2023). Wei et al. (2022) has demonstrated the emerging abilities of LLMs that only appear in larger models and cannot be predicted based on smaller models’ performance. Newer models such as GPT-4 illustrated strong capabilities in language processing and manipulation, quantitative reasoning, as well as planning and learning (Bubeck et al., 2023). Moreover, Gurnee and Tegmark (2023) has shown LLMs do have structured knowledge on both space and time using Llama-2 family models; they illustrated the model activations demonstrate linear representations of landmarks and historical events. In addition, products like ChatGPT which implement LLMs into chatbots and search engines that are easily accessible by the general public generated tremendous attention and excitement in this field (Zhao et al., 2023).
@@
Apart from their powerful ability in complex tasks, LLM has brought a new research paradigm that leverages the interaction with prompt interfaces or APIs to develop applications as opposed to trained models from scratch (Zhao et al., 2023). The performances of these LLMs to user inputs often depend on the level of prompt engineering. Common methods to improve the performance of the tasks in LLMs include in-context learning (Brown et al., 2020; Dong et al., 2023), chain-of-thought (Wei et al., 2023) and plan-and-solve (Wang et al., 2023a; Zhou et al., 2023) to elicit the step-by-step reasoning power of LLMs (Zhao et al., 2023).
@@
On the other hand, LLMs are also prone to several drawbacks that need to be considered when utilizing them. One limitation is Hallucination, where texts generated by LLMs could conflict with themselves or can not be verified (Bang et al., 2023). Another challenge of LLMs is their limit on recent knowledge since models trained on a big corpus of data with a large number of parameters typically cannot be updated in real-time (Zhao et al., 2023).
2.2 Large language model as a predictor
@@
With the exceptional performance of LLMs, studies in many fields have utilized them to liberate traditional practices. ChatGPT has been shown to achieve high accuracies for linguistic-related classification tasks, such as genre recognition, personality prediction, political opinions, and hateful speech detection (Liu et al., 2023; Kuzman et al., 2023; Amin et al., 2023; Zhang et al., 2023; Huang et al., 2023). In healthcare, ChatGPT is used to facilitate tasks such as image annotations and finding summaries (Liu et al., 2023). For example, Caruccio et al. (2024) compared the performance of ChatGPT, other pre-trained language models, and traditional machine learning models on predicting low- and medium-risk diseases based on data on symptoms and diseases. The field of law also attracts many studies in LLM predictions due to its linguistic dependence. For instance, Trautmann et al. (2022) demonstrated the natural language processing ability of the GPT models by asking it to make legal judgment predictions from long legal documents.
@@
In the field of transportation, several papers have shown the capability of LLMs to make predictions. Before the boom of LLMs, there have been several studies using pre-trained language models to predict mobility. Xue et al. (2022); Xue and Salim (2022) have been looking at converting mobility data into language descriptions, and using them as input data to fine-tune language models. With the rise of the GPT family, Wang et al. (2023b) demonstrated the powerful ability of GPTs to understand semantic sequential mobility data. In this study, each person’s historical locations, which are typically presented as structured numerical sequential data, are converted into language prompts and fed into the GPT to let it infer the next location. The predictions by LLMs outperform the traditional time-series models such as LSTM and the Multi-Head Self-Attentional (MHSA) neural network. Existing research offers promising potential for using LLMs as predictors. However, the existing literature only focuses on leveraging LLMs as a predictor in sequential mobility prediction and there has been no research exploring the ability of LLMs in travel behavior prediction tasks.
@@ 3.1.1 Large language model
Pre-training is a crucial phase in the development of LLMs, where the model is exposed to massive amounts of text data to learn the statistical properties of language, including grammar, semantics, and world knowledge. The primary objective of pre-training is to train the model to predict the next word in a sentence or sequence of words. This task, known as language modeling, allows the model to capture the statistical regularities and contextual dependencies within the training data. Note that different LLMs (e.g., BERT and GPT) may have different pre-training tasks. In this study, we introduce the pre-training based on GPT’s framework for unsupervised multitask learning (Radford et al., 2019). The objective of pre-training can be expressed as:
@@
$s^{*}=\operatorname*{arg\,max}_{s}\mathbb{P}(s\mid\text{(Input, Task)};\boldsymbol{\theta^{*}})$ (2)
where $s^{*}$ is the output sequence with the largest probability (or relatively large probability depending on the searching algorithm and degree of randomness). $s^{*}$ is generated word by word until the “[End]” token is found. For example, asking the LLM to solve an addition problem can be expressed as
[…]
@@ 3.2 Conceptual framework
Given LLMs provide a generalized multitask solver, it is possible to use them as a predictor for people’s travel behavior prediction. The “(Input, Task)” in Eq. 2 is referred to as prompt in LLMs. The performance of LLMs for many tasks highly depends on the design of the prompt. We carefully design our prompts that include 1) task descriptions, 2) travel characteristics, 3) individual attributes, and 4) guides of thinking with domain knowledge. Details will be shown in Section 3.3.
The framework of the LLM-based travel behavior prediction is shown in Figure 1. Here we use the travel mode choice task as an illustration example. The input information will be organized and embedded into the prompt (i.e., prompt design). The prompt is fed into an LLM and outputs the prediction and associated reasons. It is worth noting that no supervised training step is needed.
@@ 3.3.4 Guide of thinking with domain knowledge
The ability of LLMs to perform complex reasoning can be improved by designing prompts with strategies like Chain-of-Thought (Wei et al., 2022) and Plan-to-Solve (Wang et al., 2023a). The essential idea is to guide LLMs on how to use the given information. This is the most important part of enabling LLMs to predict travel behavior without given any training data. The prompts should include domain knowledge with human common sense. The example of travel mode choice is shown in Figure 5.
@@
The first three aspects are domain knowledge. This can be customized based on input features and the specific task. The last paragraph is used to guide LLMs for numerical comparison. It is known that arithmetic and symbolic reasoning are challenging tasks for LLMs without well-designed prompts (Rae et al., 2021). Even for a simple task like providing LLMs with three numbers A, B, and C, and asking them to sort the numbers, LLMs can make many mistakes (this problem may be mitigated with more advanced LLMs). Therefore, we add the last paragraph, telling LLMs which mode has the lowest travel time or travel cost, and how large the relative differences are. Experiments show that this tends to be an effective way to enable LLMs for arithmetic reasoning in this context.
@@ 3.3.5 Interpretation and output
The last part of the prompt is to specify the outputs of the LLMs. The outputs include two parts: 1) the predictions and 2) the reasons that support the prediction. There are two motivations to ask for the reasons. First, it gives the model interpretability and helps us to diagnose the prompt design. Moreover, asking for explanations also improves LLMs’ reasoning ability. This is similar to the Plan-to-Solve prompting (Wang et al., 2023a), where the LLM’s reasoning performance is improved when the specific calculation steps are asked to output. The example prompt for travel behavior prediction is shown in Figure 6. Note that we also specify the output format so that we can easily extract the results.

### 10 Mar 2026 v2

@@ Abstract
Travel behavior prediction is a core problem in transportation demand management and is traditionally addressed using numerical models calibrated on observed data. With recent advances in large language models (LLMs), new opportunities have emerged to model human decision-making through natural language reasoning. This study explores the use of LLMs for travel behavior prediction through two complementary frameworks. The first framework employs a zero-shot prompting strategy, where the prediction task, traveler attributes, and relevant domain knowledge are described in text, enabling the LLM to directly generate predictions without task-specific training data. The second framework uses LLM-generated text embeddings as high-level representations of travel scenarios, which are then combined with conventional supervised learning models to support prediction in small-sample settings. Empirical results show that both approaches achieve performance comparable to, and in some cases competitive with, classical models such as multinomial logit, random forest, and neural networks. These findings suggest that LLMs offer a flexible and data-efficient alternative for travel behavior prediction.
@@ 1 Introduction
Understanding human travel behavior has always been a crucial part of transportation planning. Travel behavior is a broad concept that consists of many attributes, such as the mode of transportation, the purpose of the travel, the choice of destination, and the time of departure. Consider travel mode choice as an example, which refers to how people make decisions on their means of travel. In the task of predicting mode choices, making accurate predictions is useful for both agencies and individuals. For agencies, it facilitates providing better transit services and forecasting travel conditions in the short term, while informing transportation infrastructure investments, better policies, and land use planning in the long run. In the traditional Four-Step Model for future demand and performance analysis, the mode share of a given population is estimated using mode split models [Ben-Akiva and Lerman, 1985]. For individuals, on the other hand, mode choice prediction embedded in map applications helps them to make sound travel decisions. Conventional mode choice analysis relies on using numerical data to construct a mathematical model to represent human preferences [Mo et al., 2018, 2021]. One example is the discrete choice model, where mode choice probabilities are calculated assuming people make choices to maximize their utilities [Ben-Akiva and Lerman, 1985].
@@
Recently, large language models (LLMs) have generated a tremendous amount of excitement in research due to their advanced language interaction and interpretation capabilities exhibited by-products such as ChatGPT. Having been trained on extremely large corpus and with parameters to the size of more human hundreds of billions, LLMs are able to interpret and execute diverse types of intricate tasks with a certain level of intelligence [Zhao et al., 2023]. In transportation, studies have shown the abilities of GPT in making mobility predictions which outperforms state-of-the-art time-series machine learning models [Wang et al., 2023b]. However, the question of whether LLMs can tackle the problem of travel behavior predictions is unclear.
In response to the above inquiry, this study investigates the capability of large language models (LLMs) to predict individual travel behavior, with travel mode choice and trip purpose prediction serving as a representative case studies. We propose two distinct LLM-based prediction frameworks. The first framework adopts a zero-shot prompting approach, in which carefully designed prompts describe the travel mode choice task, individual travel characteristics, and traveler attributes, while explicitly incorporating transportation domain knowledge to guide the model’s reasoning. This framework enables direct prediction through text generation without requiring any task-specific training data. The second framework leverages LLM-generated text embeddings as high-level feature representations, which are subsequently used as inputs to supervised learning models for prediction under small-sample settings. Empirical results demonstrate that both frameworks can achieve performance comparable to, and in some cases competitive with, classical travel behavior models such as multinomial logit, random forest, and neural networks, highlighting the potential of LLMs as flexible and data-efficient tools for travel behavior prediction.
@@ 2.1 Recent developments in large language models
Large Language Models (LLMs) typically refer to the language models based on the Transformer architecture with more than hundreds of billions of parameters, which have been shown to have abilities to excel in many complex tasks [Zhao et al., 2023]. Wei et al. [2022a] has demonstrated the emerging abilities of LLMs that only appear in larger models and cannot be predicted based on smaller models’ performance. Newer models such as GPT-3.5, GPT-4 [OpenAI, 2024a], Gemini [Team, 2024], and Llama [Touvron_Lavril_Izacard_Martinet_Lachaux_Lacroix_Rozière_Goyal_Hambro_Azhar_2023, Touvron et al., 2023, Teams, 2024] illustrated strong capabilities in language processing and manipulation, quantitative reasoning, as well as planning and learning [Bubeck et al., 2023]. Moreover, Gurnee and Tegmark [2023] has shown LLMs do have structured knowledge on both space and time using Llama-2 family models; they illustrated the model activations demonstrate linear representations of landmarks and historical events. In addition, products like ChatGPT which implement LLMs into chatbots and search engines that are easily accessible by the general public generated tremendous attention and excitement in this field [Zhao et al., 2023].
@@
Apart from their powerful ability in complex tasks, LLM has brought a new research paradigm that leverages the interaction with prompt interfaces or APIs to develop applications as opposed to trained models from scratch [Zhao et al., 2023]. The performances of these LLMs to user inputs often depend on the level of prompt engineering. Common methods to improve the performance of the tasks in LLMs include in-context learning, which leverages formulations of tasks and demonstrations into natural languages [Brown et al., 2020, Dong et al., 2022], chain-of-thought, which introduces intermediate reasoning [Chu et al., 2024], and plan-and-solve [Wang et al., 2023a, Zhou_Schärli_2023] to breakdown complex tasks and elicit the step-by-step reasoning power of LLMs [Zhao et al., 2023].
@@
On the other hand, apart from the fundamental abilities of LLMs in language generation, knowledge utilization, and complex reasoning, LLMs are also prone to several drawbacks that need to be considered when utilizing them [Zhao et al., 2023]. One limitation is Hallucination, where texts generated by LLMs could conflict with themselves or can not be verified [Bang et al., 2023, Huang et al., 2025]. Another challenge of LLMs is their limit on recent knowledge since models trained on a big corpus of data with a large number of parameters typically cannot be updated in real-time [Zhao et al., 2023]. A third limitation is inconsistent reasoning, where LLMs’ reasoning and answer are not always aligned [Wei et al., 2022b].
2.2 Large language model as a predictor in human mobility
@@
With the exceptional performance of LLMs, studies in many fields have utilized them to liberate traditional practices. ChatGPT has been shown to achieve high accuracies for linguistic-related classification tasks, such as genre recognition, personality prediction, political opinions, and hateful speech detection [Liu et al., 2023, Kuzman_Mozetič_Ljubešić_2023, Amin et al., 2023, Zhang et al., 2023, Huang et al., 2023]. Similarly, the utilization of LLMs in transportation also boomed in the past several year, including tasks such as forecasting in traffic, human mobility, and demand, as well as data imputation [Zhang et al., 2024].
@@
Specifically, in the realm of human mobility, a series of experiments have demonstrated the semantic and reasoning power of LLMs. Before the rise of GPT family, several studies used pre-trained language models to predict mobility. Xue et al. [2022], Xue and Salim [2022] have been looking at converting mobility data into language descriptions to fine-tune language models, while Kobayashi et al. [2023], Gong et al. , and Wu et al. [2024] examined encodings and embeddings in mobility prediction in the transformer architecture. With available commercial LLMs, Luo et al. [2024] proved GPT’s ability to infer trajectory patterns from data, while Wang et al. [2023b] initially demonstrated the powerful ability of GPTs to infer the next location with the understanding of semantic sequential mobility data and prompts converted from historical locations. The predictions by LLMs outperform the traditional time-series models such as LSTM and the Multi-Head Self-Attentional (MHSA) neural network. Additional studies illustrated how incorporating behavioral theories [Shao et al., 2024], agentic logical thinking [Wang et al., 2024a, Li et al., 2024b, a, Feng et al., 2025], pattern seeking [Qin et al., 2025, Wang et al., 2024a, Li et al., 2024b], and understanding of additional semantic data Liang et al. [2023] also facilitates the trajectory or time-series predictions.
LLMs’ planning ability has been popular in solving trip planning, one of the downstream applications of transport prediction. Leveraging multiple sources and queries [Xie et al., 2024, Fang et al., 2024, Singh et al., 2024] and breaking down into multiple steps [Tang et al., 2024, Xie and Zou, 2024] have been shown relatively successful.
Despite existing research offers promising potential for using LLMs as predictors, the existing literature only focuses on leveraging LLMs as a predictor in sequential mobility prediction and there has been very few research exploring the ability of LLMs in travel behavior prediction tasks. In terms of travel choice prediction, a few studies analyzed how to incorporate historical personalized information to improve prediction [Zhai et al., 2024, Wang et al., 2024b, Chen et al., 2024]. Liu et al. [2024] introduced persona loading and few-shots to improve the misalignment of LLM predictions and the human behavior. However, there hasn’t been a systematic study on understanding the zero-shot learning in travel behavior predictions, especially in terms of the tasks of mode choice and travel purpose inference with LLM, which we will focus in this study.
@@ 3.1 Preliminaries
Pre-training is a crucial phase in the development of LLMs, where the model is exposed to massive amounts of text data to learn the statistical properties of language, including grammar, semantics, and world knowledge. The primary objective of pre-training is to train the model to predict the next word in a sentence or sequence of words. This task, known as language modeling, allows the model to capture the statistical regularities and contextual dependencies within the training data. Note that different LLMs (e.g., BERT and GPT) may have different pre-training tasks. In this study, we introduce the pre-training based on GPT’s framework for unsupervised multitask learning [Radford et al., 2019]. The objective of pre-training can be expressed as:
@@
$\boldsymbol{s}^{*}=\operatorname*{arg\,max}_{\boldsymbol{s}}\mathbb{P}(\boldsymbol{s}\mid\text{(Input, Task)};\boldsymbol{\theta^{*}})$ (2)
where $\boldsymbol{s}^{*}$ is the output sequence with the largest probability (or relatively large probability depending on the searching algorithm and degree of randomness). $\boldsymbol{s}^{*}$ is generated word by word until the “[End]” token is found. For example, asking the LLM to solve an addition problem can be expressed as
$\text{``The answer is 28 [End]''}=\operatorname*{arg\,max}_{\boldsymbol{s}}\mathbb{P}(\boldsymbol{s}\mid\text{``What is 3+25? [End]''};\boldsymbol{\theta^{*}})$ (3)
@@
Besides generating text sequences, LLMs can also be used to obtain text embeddings using the internal layers before the final text output layer. Specifically, denote the LLM’s with embedding outputs as $f(\cdot;\boldsymbol{\theta}^{*})$. Then the text embedding (denoted as $\boldsymbol{h}$) for the input task can be obtained as:
$\boldsymbol{h}=f\left(\text{(Input, Task)};\boldsymbol{\theta}^{*}\right)$ (4)
$\boldsymbol{h}\in\mathbb{R}^{H}$ is usually use as features for many downstream supervised learning tasks as the extracted information of text. Its dimension $H$ depends on specific LLM architectures. Given the generalizability of LLMs, $\boldsymbol{h}$ could potentially increase the prediction performance of models with small sample sizes.
@@ 3.2 Conceptual framework
Given LLMs provide a generalized multitask solver, it is possible to use them as a predictor for people’s travel behavior prediction. In this study, we propose two different frameworks to incorporate LLM.
The “(Input, Task)” in Eq. 2 is referred to as prompt in LLMs. The performance of LLMs for many tasks highly depends on the design of the prompt. We carefully design our prompts that include 1) task descriptions, 2) structural data, 3) descriptive data, and 4) guides for thinking with domain knowledge. Details will be shown in Section 3.3.
The frameworks of the LLM-based travel behavior prediction is shown in Figure 1. Here, we use a travel mode choice task and a travel purpose prediction task as the illustration examples. The input information will be organized and embedded into the prompt (i.e., prompt design). The prompt is fed into an LLM and outputs the prediction and associated reasons. It is worth noting that no supervised training step is needed.
@@
After processing of LLM, we may get two different outputs depending on which prediction framework we use. For direct prediction based on prompts, we will use text-generation LLM (e.g., GPT 3.5) and ask the model to generate the predicted results and reasons. In terms of using prompt embedding as features, we will get a hidden vector $\boldsymbol{h}_{n}$ from the LLM (e.g., text-embedding model [OpenAI, 2024b] in the Open AI platform) for sample $n$. Then we train a supervised learning model (e.g., multinomial logit) using $\boldsymbol{h}_{n}$ as features, and use the trained model to predict the travel behavior.
@@ 3.3.4 Guide of thinking with domain knowledge
The ability of LLMs to perform complex reasoning can be improved by designing prompts with strategies like Chain-of-Thought [Wei et al., 2022a] and Plan-to-Solve [Wang et al., 2023a]. The essential idea is to guide LLMs on how to use the given information. This is the most important part of enabling LLMs to predict travel behavior without given any training data. The prompts should include domain knowledge with human common sense. The example of travel mode choice is shown in Figure 5.
@@
The first three aspects are domain knowledge. This can be customized based on input features and the specific task. The last paragraph is used to guide LLMs for numerical comparison. It is known that arithmetic and symbolic reasoning are challenging tasks for LLMs without well-designed prompts [Rae et al., 2021]. Even for a simple task like providing LLMs with three numbers A, B, and C, and asking them to sort the numbers, LLMs can make many mistakes (this problem may be mitigated with more advanced LLMs). Therefore, we add the last paragraph, telling LLMs which mode has the lowest travel time or travel cost, and how large the relative differences are. Experiments show that this tends to be an effective way to enable LLMs for arithmetic reasoning in this context.
@@ 3.3.5 Interpretation and output
The last part of the prompt is to specify the outputs of the LLMs. The outputs include two parts: 1) the predictions and 2) the reasons that support the prediction. There are two motivations to ask for the reasons. First, it gives the model interpretability and helps us to diagnose the prompt design. Moreover, asking for explanations also improves LLMs’ reasoning ability. This is similar to the Plan-to-Solve prompting [Wang et al., 2023a], where the LLM’s reasoning performance is improved when the specific calculation steps are asked to output. The example prompt for travel behavior prediction is shown in Figure 6. Note that we also specify the output format so that we can easily extract the results.
@@ 3.3.6 Summary
3.4 Embedding as features
Large Language Models (LLMs) encode semantic and contextual information into high-dimensional embedding vectors that can be leveraged for downstream prediction tasks. In the context of travel behavior prediction, we extract an embedding $\boldsymbol{h}_{n}$ for each sample nn—typically representing a traveler, trip record, or textual description of the travel context. This embedding $\boldsymbol{h}_{n}$ serves as a compact yet expressive representation of the traveler’s behavioral features, preferences, and context.
To predict the travel behavior label $y_{n}$ (e.g., mode choice, trip purpose, or destination), we use $\boldsymbol{h}_{n}$ as input to a supervised learning model parameterized by $\boldsymbol{\beta}$. The model is trained to minimize the empirical loss over the training dataset $\mathcal{D}^{\text{Train}}$:
$\boldsymbol{\beta}^{*}=\arg\min_{\boldsymbol{\beta}}\sum_{n\in\mathcal{D}^{\text{Train}}}\ell_{n}\left(\text{SupervisedModel}(\boldsymbol{h}_{n};\boldsymbol{\beta}),y_{n}\right)$ (5)
Here, $\ell_{n}(\cdot,\cdot)$ denotes a sample-wise loss function that quantifies the discrepancy between the predicted and true labels. For classification tasks, $\ell_{n}$ is typically the cross-entropy loss, while for regression tasks, it may be the mean squared error or other appropriate metrics.
This approach decouples representation learning (handled by the LLM) from the prediction model, allowing flexibility in choosing downstream models—such as logistic regression, random forests, gradient boosting machines, or neural networks—based on task-specific requirements and computational considerations. Notably, the use of $\boldsymbol{h}_{n}$ as input enables generalization across diverse travel contexts, especially when the embeddings capture high-level behavioral semantics informed by large-scale pretraining.